# Skill-informed Data-driven Haptic Nudges for High-dimensional Human Motor Learning

### 13 Mar 2026 v1

@@ 3.1 Behavioral Model
We model the human motor learning behavior as an Input-Output Hidden Markov Model (IOHMM), as illustrated in Fig. 2222The Python package by Mogeng et al. that was used to build and train our IOHMM models can be found at https://github.com/Mogeng/IOHMM.git. The hidden state $h_{k}\in\{1,\dots,N\}$ represents one of $N$ discrete, latent motor skill states at trial $k$. To maintain a reasonable number of parameters, the input covariates vector $\mathbf{x}_{k}$ is composed of two experimental variables: the task complexity, captured by the slope angle $u_{k}$ of the target pair, and the haptic guidance, represented by the finger index $a_{k}$ to which the haptic feedback is applied (an index from $0\text{--}5$, where 0 denotes no nudge). This prescribed haptic feedback consisted of a tactile stimulus: two sequential vibration bursts, each lasting 150150~ms, separated by 20002000~ms, delivered to the fingertip indexed $a_{k}$. The strength of this nudge is chosen such that it is clearly perceived by the participant, but doesn’t impact the finger motion. A target pair consists of targets for trial $k$ and k1k-1, with their numbering convention following $\{0:(\mathcal{T}(2),\mathcal{T}(3)),\ 1:(\mathcal{T}(1),\ \mathcal{T}(2)),\ 2:(\mathcal{T}(1),\mathcal{T}(3)),\ 3:(\mathcal{T}(1),\mathcal{T}(4)),\ 4:(\mathcal{T}(2),\mathcal{T}(4)),\ 5:(\mathcal{T}(3),\mathcal{T}(4))\}$.
The evolution of the latent motor skill state from trial $k$ to $k+1$ is governed by the transition probability distribution $\mathbb{P}(h_{k+1}|h_{k},\mathbf{x}_{k})$. The observed emissions (outputs) $\mathbf{o}_{k}=[\texttt{RE}_{k},\texttt{SoT}_{k}]^{\top}$ represent the performance metrics of the learner: Reaching Error (RE) and Straightness of Trajectory (SoT). The distribution of these outputs is conditional on the current state and the input covariates, defined by the emission models $\mathbb{P}(\mathbf{o}_{k}|h_{k},\mathbf{x}_{k})$. Together with the initial state distribution $\mathbb{P}(h_{0})$, these components form a predictive model of motor learning behavior under haptic guidance as an IOHMM.
@@ 3.2 Estimating the Model Parameters
The parameters of the IOHMM were estimated using the gameplay data collected from the first five participants in the Heuristic Group playing the target capture game with heuristic nudging policy (see Sec. 4.1, and 5.1). Unlike standard HMMs, which utilize fixed probability tables, the IOHMM employs parameterized functions to map the input covariates $\mathbf{x}_{k}$ to state transitions and emission distributions.
@@ 3.2.1 Initial and Transition Models
The initial state distribution $\mathbb{P}(h_{0}|\mathbf{x}_{0})$ and the state transition probabilities $\mathbb{P}(h_{k+1}|h_{k},\mathbf{x}_{k})$ were modeled using Multinomial Logistic Regression (MNL). Formally, the probability of transitioning from a state $h_{k}=i$ to a state $h_{k+1}=j$ given input $\mathbf{x}_{k}$ is defined by the softmax function
@@
where $\mathbf{w}_{ij}$ and $b_{ij}$ are the coefficient vector and intercept term associated with the transition $i\to j$. To prevent overfitting given the limited training data, we applied $L_{2}$ regularization (Ridge regression) during the fitting process. The initial state distribution follows a randomized initialization since we do not enforce a semantic ordering of the latent states, and only perform a post-hoc re-ordering for visualization and interpretability.
3.2.2 Emission Model
The continuous emission outputs $\mathbf{o}_{k}$ (comprising RE and SoT) were modeled using a linear Gaussian function. For a given state $h_{k}=i$, the observation is assumed to be drawn from a multivariate normal distribution where the mean is a linear function of the inputs
@@ 3.2.2 Emission Model
where $\mathbf{V}_{i}$ is the coefficient matrix, $\mathbf{c}_{i}$ is the intercept vector, and $\mathbf{\Sigma}_{i}$ is the covariance matrix for state $i$. The parameters for the emission means were estimated using Ordinary Least Squares (OLS) regression. To enhance model generalization, we applied Elastic-Net regularization, which linearly combines $L_{1}$ and $L_{2}$ penalties, promoting sparsity while maintaining stability.
@@ 3.2.3 Training Algorithm
Since the dependencies in an IOHMM are parametric, the standard Baum-Welch algorithm cannot be applied directly. Instead, we utilized a Generalized Expectation-Maximization (GEM) algorithm. In the Expectation (E) step, the forward-backward algorithm computes the posterior probabilities of being in state $i$ at trial $k$, denoted $\gamma_{k}(i)$, and the joint probabilities of transitioning from $i$ to $j$, denoted $\xi_{k}(i,j)$. In the Maximization (M) step, rather than counting frequencies, these posterior probabilities serve as sample weights for the supervised learning tasks. The transition parameters ($\mathbf{w},b$) are updated by fitting weighted MNL classifiers, and the emission parameters ($\mathbf{V},\mathbf{c},\mathbf{\Sigma}$) are updated by fitting weighted linear regressors.
@@ 3.2.4 Model Selection
Figure 3: Estimated Behavioral Model: (a) Predicted RE and SoT means out of the emission model for the ordered skill states, and (b) State Transition Matrices extracted from the estimated human motor learning behavioral model show how participants tend to transition from low latent skill states to high latent skill states under the effect of the inputs $u_{k}=-1.57,-1.11$ and $a_{k}={0,\cdots,5}$. The latent skill states are ordered best-to-worst (06)(0-6) based on their mean output RE emission values. The $\texttt{nudge}=0$ value represents no nudging.
@@ 3.3 State Transition Matrices
To study how the latent skill states probabilistically transition as a function of target pair and nudge feedback, we plot the transition function as state transition matrices for the pair of inputs $\mathbf{x}_{k}$. For interpretability, we first order the latent motor skill states from best $(0)$ to worst $(6)$ based on the increasing value of the average of predicted $\text{mean}(\texttt{RE})+\text{mean}(\texttt{SoT})$ out of the emission model (Fig. 3), where the $\text{mean}(\cdot)$ is taken over all inputs $\mathbf{x}_{k}$. We then analyze these transitions for the hardest and easiest target pair, denoted by their slopes, under the effect of different nudge feedbacks, where 0 refers to no nudge. The order of difficulty of various target pairs is obtained through studying the number of gameplay trials it takes for SoT to converge to a baseline value of $0.3$ on average across the five participants whose data is used to estimate the IOHMM. Target pair 0 (see Section 3.1 for target pair numbering convention) with slope $\{\pm 1.57\}$ is the easiest with participants converging to an SoT of $0.3$ by trial $19$ on average, whereas target pair $2$ with slope angle $\{-1.11,2.03\}$ is the hardest, with SoT converging to $0.3$ by trial $210$ on average. Fig. 3 shows the state transition matrices corresponding to the target pair 0.
The probabilities show how participants tend to transition from low-performing motor skill states to high-performing motor skill states, ideally imparting an expected (albeit imperfect) lower triangular structure to these matrices owing to the new state ordering.
The figure also reveals that for the target pair 0 with slope angle 1.57-1.57, the participants tend to stay in high performing skill states 0,1{0,1} without needing any nudge. Additionally, a nudge to finger $1$, which is the thumb, helps them transition from lower to higher performing skill states more prominently. This was also observed in the experiments, where the thumb abduction-adduction majorly led to the screen cursor’s vertical motion. Furthermore, Fig. 3 also shows that for the target pair $2$ with slope angle 1.11-1.11, nudges are more helpful for transitioning from lower to higher performing skill states compared to nudges for target pair 0. In both these cases, finger $5$, which corresponds to the pinky, had the least effect in transitioning away from lower performing skill states. This is also observed in the experiments, in that the pinky did not contribute much to the cursor motion owing to its relatively lower range of motion.
@@ 4 Skill-based Nudging Prescription Design
The application of haptic force feedback offers a promising avenue for expediting motor skill acquisition by providing physical cues that guide the learner toward desired behaviors. Nevertheless, the efficacy of such interventions in high-dimensional spaces is constrained by two primary obstacles. First, true motor proficiency is a latent variable, an internal state of the neural controller that is inaccessible for direct measurement [12]. Conventional feedback strategies often depend on observable, low-dimensional error metrics as proxies for skill. However, strictly minimizing these kinematic errors does not ensure the development of robust, high-dimensional coordination strategies and may inadvertently reinforce sub-optimal compensatory habits. Second, in contrast to low-dimensional scenarios where reference trajectories are well-defined and can be tracked directly [esmatloo2020fingertip, sullivan2021haptic], complex high-degree-of-freedom tasks lack a unique “ground truth” trajectory. This redundancy makes it non-trivial to design intuitive feedback that effectively resolves the learner’s kinematic ambiguity. To address these limitations, we propose a paradigm shift from performance-error correction to skill-state-aware guidance. By utilizing the probabilistic model of human motor learning estimated from input-output data (detailed in Section 3), our approach synthesizes personalized haptic nudges. These interventions are designed to regulate the learner’s exploration of the control manifold, thereby fostering the discovery of efficient coordination patterns and accelerating skill learning.
@@ 4.1 Heuristic Nudging Policy
To establish a performance baseline, we synthesized a heuristic feedback policy informed by pilot studies. This strategy aims to assist the user by incrementally refining their hand configuration to achieve the current objective. For a specific target $t_{k}$ presented in trial $k$, we first calculate a theoretical optimal posture, $\theta^{\textup{opt}}_{k}$, by applying the pseudo-inverse of the BoMI mapping matrix, such that $\theta^{\textup{opt}}_{k}=C^{\dagger}t_{k}$. We then evaluate the difference between this ideal configuration and the participant’s actual hand posture, $\theta^{\textup{curr}}_{k}$, at the onset of the trial. The deviation of each finger $i$ is quanepsied in joint space as the Sum of Squared Errors (SSE) between the current and optimal angles as
@@
The specific finger to receive the haptic nudge is sampled stochastically from this distribution. The underlying logic is that minimizing $d_{i}$ idenepsies the finger capable of producing the most significant cursor position with the least kinematic effort. Consequently, this heuristic functions as a local refinement mechanism. It serves as a computationally lightweight standard against which we evaluate our optimal, POMDP-driven nudging policy.
@@ 4.2 POMDP-based Optimal Nudging Policy
To synthesize an active intervention strategy, we recast the predictive IOHMM framework (Section 3) into a control-theoretic POMDP. The primary objective is to compute a policy $\pi$ that maps a probabilistic belief of the user’s hidden motor state to specific haptic nudge feedback. This policy is optimized to minimize the cumulative expected deviation in performance metrics (RE and SoT) over a long horizon. The POMDP is characterized by the tuple $\langle\mathcal{S,A},T,R,\Omega,O\rangle$, with components adapted from the underlying IOHMM structure.
@@
Transition and Emission Models: The system dynamics are decomposed into the evolution of the human’s skill and the stochastic progression of the game targets. Given that targets are drawn uniformly randomly (excluding the immediate predecessor) and independent of the motor state, the composite transition model is written as
$T=\mathbb{P}(s_{k+1}|s_{k},a_{k})=\mathbb{P}(h_{k+1}|h_{k},u_{k},a_{k})\times\mathbb{P}(t_{k}).$ (6)
Similarly, the emission model is inherited directly from the IOHMM structure $O=\mathbb{P}(o_{k}|s_{k},a_{k})=\mathbb{P}(o_{k}|h_{k},u_{k},a_{k})$. It is important to note that $u_{k}$ is deterministically determined from the target pair $\{t_{k-1},t_{k}\}$.
@@
$R(s_{k},a_{k})=-(\mathbb{E}[C_{\textup{imm}}(s_{k},a_{k})]+w_{\textup{g}}\mathbb{E}[C_{\textup{gen}}(h_{k+1},s_{k},a_{k})]),$ (7)
where $\mathbb{E}[C_{\textup{imm}}(s_{k},a_{k})]$ denotes the expected cost for the current trial, while $\mathbb{E}[C_{\textup{gen}}(h_{k+1},s_{k},a_{k})]$ evaluates the expected performance on all possible tasks at the next latent state $h_{k+1}$. The weighting parameter was fixed at $w_{\textup{g}}=1$. The immediate performance penalty is a linear combination of the predicted means for RE and SoT as
$\mathbb{E}[C_{\textup{imm}}(s_{k},a_{k})]=\mu_{\textup{RE}}(h_{k},u_{k},a_{k})+w_{\textup{SoT}}\mu_{\textup{SoT}}(h_{k},u_{k},a_{k}),$ (8)
where $\mu_{\textup{RE}}$ and $\mu_{\textup{SoT}}$ are the predicted RE and SoT means from the emission model, and $w_{\textup{SoT}}=2$ scales the relative importance of SoT. The expected future utility is calculated by marginalizing over the transition probabilities provided by the IOHMM as
$\mathbb{E}[C_{\textup{gen}}(h_{k+1},s_{k},a_{k})])=\sum_{h_{k+1}}\mathbb{P}(h_{k+1}|s_{k},a_{k})\boldsymbol{\cdot}C_{\textup{gen}}(h_{k+1}),$ (9)
@@ 4.3 Approximate POMDP Solution via QMDP
The state space of the POMDP is characterized by the belief space, a continuous, high-dimensional simplex representing the set of all valid probability distributions over the underlying POMDP states $s_{k}\in\mathcal{S}$. Given the computational intractability of applying standard infinite-horizon value iteration directly to this continuous space, we adopt the QMDP approximation method [littman1995learning]. This algorithm computes $Q_{MDP}(s,a)$ values for the underlying MDP by utilizing only the transition and reward models, effectively bypassing the observation model to render the problem tractable. Although this approach technically assumes full state observability, an assumption that does not hold in our partial information setting, it yields a robust policy $\pi^{*}$ that can be applied to states sampled from the current belief distribution. To further robusepsy the policy against model imperfections stemming from limited training data, the MDP is solved via infinite-horizon soft-value iteration. The value update rule follows the soft Bellman optimality equation $V(s)\leftarrow\alpha\log\sum_{a\in\mathcal{A}}\exp\left(\frac{Q(s,a)}{\alpha}\right)$, where $\alpha$ denotes the temperature parameter. The corresponding action-value function $Q(s,a)$ is defined as
$Q(s_{k},a_{k})=R(s_{k},a_{k})+\gamma\sum_{s_{k+1}\in\mathcal{S}}\mathbb{P}(s_{k+1}|s_{k},a_{k})V(s_{k+1}).$ (11)
The algorithm iterates until the value function $V(s_{k})$ converges across the state space. The final output is a stochastic policy $\pi(a|s_{k})$, formulated as a softmax distribution over the $Q$-values, which assigns a selection probability to each haptic nudge prescription. For this implementation, we used $\alpha=0.2$ and $\gamma=0.98$.
Real-Time Policy Execution and Belief Update
The derived stationary policy is integrated into the experiment for real-time execution. Throughout the experiment, the system maintains a running belief state $b_{k}:=b(s_{k})$, representing the posterior probability distribution $\mathbb{P}(s_{k}|o_{k},a_{k},s_{k-1})$. The selection of the specific finger index for haptic feedback proceeds in two stages: first, a state hypothesis $s_{k}$ is sampled from the current belief distribution $s_{k}\sim b_{k}$; second, the action $a_{k}$ is sampled from the policy distribution $\pi^{*}(a|s_{k})$ corresponding to that sampled state. Following the prescription of the haptic nudge by the BoMI, the system records the participant’s trial performance metrics $o_{k}=\{\texttt{RE}_{k},\texttt{SoT}_{k}\}$. The belief state is then advanced to the subsequent time step via the Bayesian filter
@@
where $\eta$ is the normalization constant. This updated belief $b_{k+1}$ encapsulates the new system knowledge and is used in the decision process for the next trial.
Figure 4: Task Performance: (a) Mean SoT curves (with $95\%$ confidence intervals) of the three group participants, and (b) the mean number of trials (along with the scatter plot taken across participants) to converge to a specific SoT threshold under the three nudging policies shows fastest convergence for participants trained on POMDP-based optimal nudging policy, followed by the heuristic policy. The low p-values from two-tailed tests from the linear mixed model fits show a decrease in SoT across trials and in the number of trials required by the POMDP group for various SoT thresholds. Each scatter plot point represents the number of trials taken by a participant to achieve the SoT threshold.

### 13 Apr 2026 v2

@@ 3.1 Behavioral Model
We model the human motor learning behavior as the IOHMM [24] illustrated in Fig. 2. The hidden state $h_{k}\in\{1,\dots,N\}$ represents one of $N$ discrete, latent motor skill states at trial $k$. To maintain a reasonable number of parameters, the input vector $\mathbf{x}_{k}$ is composed of two experimental variables: the task type, captured by the slope angle $u_{k}$ of the target pair, and the haptic guidance, represented by the finger index $a_{k}$ to which the haptic feedback is applied (an index from $0\text{--}5$, where 0 denotes no nudge, and 151-5 denote fingers from thumb to the pinky). A target pair consists of targets for trial $k$ and k1k-1, with their numbering convention following $\{0:(\mathcal{T}_{2},\mathcal{T}_{3}),\ 1:(\mathcal{T}_{1},\mathcal{T}_{2}),\ 2:(\mathcal{T}_{1},\mathcal{T}_{3}),\ 3:(\mathcal{T}_{1},\mathcal{T}_{4}),\ 4:(\mathcal{T}_{2},\mathcal{T}_{4}),\ 5:(\mathcal{T}_{3},\mathcal{T}_{4})\}$.
The evolution of the latent motor skill state from trial $k$ to $k+1$ is governed by the transition probability distribution $\mathbb{P}(h_{k+1}|h_{k},\mathbf{x}_{k})$. The outputs $\mathbf{o}_{k}=[\texttt{RE}_{k},\texttt{SoT}_{k}]^{\top}$ represent the performance metrics of the learner: Reaching Error (RE) and Straightness of Trajectory (SoT). The distribution of these outputs is conditional on the current state and the inputs, defined by the conditional distribution $\mathbb{P}(\mathbf{o}_{k}|h_{k},\mathbf{x}_{k})$. Together with the initial state distribution $\mathbb{P}(h_{0})$, these components constitute the IOHMM of motor learning behavior under haptic guidance.
@@ 3.2 Estimating the Model Parameters
The parameters of the IOHMM were estimated using the experimental data collected from the first five participants in the Heuristic Group performing the target capture task with heuristic nudging policy (see Section 4.1, and 5.1). For data-efficiency, functional approximation is used to estimate the state-transition and output probabilities.
@@ 3.2.1 Initial and Transition Models
The initial state distribution $\mathbb{P}(h_{0}|\mathbf{x}_{0})$ and the state transition probabilities $\mathbb{P}(h_{k+1}|h_{k},\mathbf{x}_{k})$ were modeled using Multinomial Logistic Regression (MNL) [12]. Formally, the probability of transitioning from a state $h_{k}=i$ to a state $h_{k+1}=j$ given input $\mathbf{x}_{k}$ is defined by the softmax function
@@
where $\mathbf{w}_{ij}$ and $b_{ij}$ are the coefficient vector and intercept term associated with the transition $i\to j$. To prevent overfitting given the limited training data, we applied $L_{2}$ regularization (Ridge regression) [11] during the fitting process. The initial state distribution follows a randomized initialization since we do not enforce a semantic ordering of the latent states, and only perform a post-hoc re-ordering for visualization and interpretability.
3.2.2 Output Model
The outputs $\mathbf{o}_{k}$ (comprising RE and SoT) were modeled using a linear Gaussian function. For a given state $h_{k}=i$ and input $\mathbf{x}_{k}$, the observation is assumed to be drawn from a multivariate normal distribution where the mean is a linear function of the inputs
@@ 3.2.2 Output Model
where $\mathbf{V}_{i}$ is the coefficient matrix, $\mathbf{c}_{i}$ is the intercept vector, and $\mathbf{\Sigma}_{i}$ is the covariance matrix for state $i$. To estimate the model parameters, the associated loss in the overall loss function is chosen as Ordinary Least Squares (OLS). To enhance model generalization, we applied Elastic-Net regularization [27], which linearly combines $L_{1}$ and $L_{2}$ penalties, promoting sparsity while maintaining stability.
@@ 3.2.3 Training Algorithm
Since the dependencies in an IOHMM are parametric, the standard Baum-Welch algorithm cannot be applied directly. Instead, we utilized the Generalized Expectation-Maximization algorithm implemented in [bengio1996input].
@@ 3.2.4 Model Selection
Figure 3: Estimated Behavioral Model: (a) Predicted RE and SoT means out of the emission model for the ordered skill states, and (b) State Transition Matrices extracted from the estimated human motor learning behavioral model show how participants tend to transition from low latent skill states to high latent skill states under the effect of the input slope angles $u_{k}=-1.57,-1.11$ (corresponding to target pairs $\{0,2\}$) and the nudged finger indices $a_{k}={0,\cdots,5}$. The latent skill states are ordered best-to-worst (06)(0-6) based on their mean output RE emission values. The $\texttt{nudge}=0$ value represents no nudging.
@@ 3.3 State Transition Matrices
We now analyze the motor learning behavior of the participants subject to task type and the nudge feedback captured by the estimated IOHMM model. Accordingly, to study how their latent skill states probabilistically transition as a function of target pair and nudge feedback, we plot the transition function as state transition matrices for the pair of inputs $\mathbf{x}_{k}$. For interpretability, we first order the latent motor skill states from best $(0)$ to worst $(6)$ based on the increasing value of the average of predicted $(\text{mean}(\texttt{RE})+\text{mean}(\texttt{SoT}))$ outputs for all inputs (Fig. 3). We then analyze these transitions for the hardest and easiest target pair, denoted by their slopes, under the effect of different nudge feedbacks, where 0 refers to no nudge. The order of difficulty of various target pairs is obtained through studying the number of experiment trials it takes for SoT to converge to a baseline value of $0.3$ on average across the five participants whose data is used to estimate the IOHMM. Target pair $0:(\mathcal{T}_{2},\mathcal{T}_{3})$ with slope angle $\{\pm 1.57\}$ is the easiest with participants converging to an SoT of $0.3$ by trial $19$ on average, whereas target pair $2:(\mathcal{T}_{1},\mathcal{T}_{3})$ with slope angle $\{-1.11,2.03\}$ is the hardest, with SoT converging to $0.3$ by trial $210$ on average. Fig. 3 shows the state transition matrices corresponding to the target pair 0 and $2$.
The probabilities show how participants tend to transition from low-performing motor skill states to high-performing motor skill states, imparting a lower triangular dominant structure to these matrices owing to the mean performance-based skill ordering.
The figure also reveals that for both the target pairs, the participants tend to stay in high performing skill states $\{0,1\}$ without needing any nudge. Additionally, for the target pair 0, a nudge to finger $1$, which is the thumb, helps them transition from lower to higher performing skill states more prominently. This was also observed in the experiments, where the thumb abduction-adduction majorly led to the screen cursor’s vertical motion. Furthermore, Fig. 3 also shows that for the target pair $2$, nudges are more helpful for transitioning from lower to higher performing skill states compared to nudges for target pair 0. It should be noted that the observed nudging effect with respect to different finger indices are dependent on the forward BoMI mapping matrix $C$ constructed via the consistent calibration performed for all participants.
@@ 4 Skill-based Nudging Prescription Design
Building on the premise that true motor skill is latent and high-dimensional tasks lack unique reference trajectories, we shift from conventional performance-error correction to skill-state-aware guidance. In this section, we first outline a heuristic nudging policy aimed at local kinematic refinement. Subsequently, leveraging the probabilistic behavioral model of motor learning (detailed in Section 3), we formulate a POMDP to synthesize optimal, personalized haptic nudges designed to regulate motor exploration, resolve kinematic ambiguity, and accelerate the discovery of efficient coordination strategies.
@@ 4.1 Heuristic Nudging Policy
We synthesized a heuristic feedback policy informed by pilot studies. This strategy aims to assist the user by incrementally refining their hand configuration to achieve the current objective by nudging the finger capable of the largest performance error reduction with the least change in finger pose. For a specific target $t_{k}$ presented in trial $k$, we first calculate a theoretical optimal posture, $\theta^{\textup{opt}}_{k}$, by applying the pseudo-inverse of the BoMI mapping matrix, such that $\theta^{\textup{opt}}_{k}=C^{\dagger}t_{k}$. We then evaluate the difference between this ideal configuration and the participant’s actual hand posture, $\theta^{\textup{curr}}_{k}$, at the onset of the trial. The deviation of each finger $i$ is quantified in joint space as the Sum of Squared Errors (SSE) between the current and optimal angles as
@@
The specific finger to receive the haptic nudge is sampled from this distribution. Here, minimizing $d_{i}$ identifies the finger capable of producing the most significant cursor position with the least kinematic effort. Consequently, this heuristic functions as a local refinement mechanism. It serves as a computationally lightweight standard against which we evaluate our optimal, POMDP-driven nudging policy.
@@ 4.2 POMDP-based Optimal Nudging Policy
To synthesize an active intervention strategy, we treat the nudge input in the IOHMM as a control variable and design a control policy using a POMDP framework. This policy $\pi$ maps the posterior distribution over human motor skill state (called belief) to specific haptic nudge feedback.
@@
Transition and Output Models: Given that at target $t_{k-1}$, the next target $t_{k}$ is drawn uniformly randomly from $\mathcal{T}\backslash t_{k-1}$ and is independent of the motor state, the composite transition model is written as
$\mathbb{P}(s_{k+1}|s_{k},a_{k})=\mathbb{P}(h_{k+1}|h_{k},u_{k},a_{k})\times\mathbb{P}(t_{k}|t_{k-1}).$ (6)
Similarly, the output model $\mathbb{P}(o_{k}|h_{k},u_{k},a_{k})$ is inherited directly from the IOHMM structure It is important to note that $u_{k}$ is completely determined from the target pair $\{t_{k-1},t_{k}\}$.
@@
$R(s_{k},a_{k})=$- $\left(C_{\textup{imm}}(h_{k},u_{k},a_{k})\right$.
$\left.+w_{\textup{g}}C_{\textup{gen}}(h_{k+1},h_{k},u_{k},a_{k})\right),$ (7)
where $C_{\textup{imm}}(s_{k},a_{k})$ denotes the cost for the current trial, while $C_{\textup{gen}}(h_{k+1},s_{k},a_{k})$ evaluates the expected performance of all possible inputs over all possible next latent states $h_{k+1}$. The weighting parameter was fixed at $w_{\textup{g}}=1$. The immediate performance penalty is a linear combination of the predicted means for RE and SoT as
$C_{\textup{imm}}=\mu_{\textup{{RE}}}(h_{k},u_{k},a_{k})+w_{\textup{{SoT}}}\mu_{\textup{{SoT}}}(h_{k},u_{k},a_{k}),$ (8)
where $\mu_{\textup{{RE}}}$ and $\mu_{\textup{{SoT}}}$ are the predicted RE and SoT means from the output model, and $w_{\textup{{SoT}}}=2$ scales the relative importance of SoT. The second term in the reward $C_{\textup{gen}}$ captures the versatility of the skill state at stage $k+1$, and is defined by
$C_{\textup{gen}}=\mathbb{E}_{h_{k+1}}[\mathbb{P}(h_{k+1}|h_{k},u_{k},a_{k})\boldsymbol{\cdot}C_{\textup{g}}(h_{k+1})],$ (9)
@@ 4.3 Approximate POMDP Solution via QMDP
A POMDP can be interpreted as an MDP in the belief space, a continuous, high-dimensional simplex representing the set of all probability distributions over the underlying POMDP states $s_{k}\in\mathcal{S}$. The belief state $b_{k}:=b(s_{k})$ represents the posterior probability distribution $\mathbb{P}(s_{k}|o_{k},a_{k},s_{k-1})$ and is iteratively computed given observation $o_{k}$ as
@@
where $\eta$ is the normalization constant. Given the computational intractability of applying standard infinite-horizon value iteration directly to this continuous belief space, we adopt the QMDP approximation method [littman1995learning] to obtain a near-optimal nudging policy. QMDP simplifies the problem by solving the underlying MDP, under the assumption of full state observability, to obtain the optimal action-value function $Q_{\textup{MDP}}(s_{k},a_{k})$ by iteratively solving
$Q_{\textup{MDP}}(s_{k},a_{k})$ $=R(s_{k},a_{k})+\gamma\sum_{s_{k+1}\in\mathcal{S}}\mathbb{P}(s_{k+1}|s_{k},a_{k})V(s_{k+1}),$
$V(s_{k})$ $=\max_{a\in\mathcal{A}}Q_{\textup{MDP}}(s_{k},a),$ (12)
until the value function $V(s_{k})$ converges across the state space. The QMDP algorithm then samples a state from the posterior belief and applies the associated optimal action. To further robustify the policy against model imperfections stemming from limited training data, the MDP is solved via infinite-horizon soft-value iteration [4]. The value update rule follows the soft Bellman update wherein the maximum operator in Bellman equation (4.3) is replaced with a softmax operator with parameter $\alpha$ as
$V(s_{k})\leftarrow\alpha\log\sum_{a\in\mathcal{A}}\exp\left(\frac{Q_{\textup{MDP}}(s_{k},a)}{\alpha}\right).$ (13)
Small values of $\alpha$ correspond to near optimality, while larger values correspond to a random action selection. This parameter acts as a regularizer and guards against model imperfections. The final output is a stochastic policy $\pi^{*}(a|s_{k})$, where action $a$ is selected with probability
$p(Q_{\textup{MDP}}(s_{k},a))\propto\frac{\exp\left(Q_{\textup{MDP}}(s_{k},a)/\alpha\right)}{\sum_{a^{\prime}}\exp\left(Q_{\textup{MDP}}(s_{k},a^{\prime})/\alpha\right)}.$ (14)
For this implementation, we used $\alpha=0.2$ and $\gamma=0.98$. These values were obtained through Monte Carlo simulations of the experiment, using the trained IOHMM as a generative model to maximize the mean policy performance.
Real-Time Policy Execution
The derived stationary policy is integrated into the experiment for real-time execution. Throughout the experiment, the system maintains a running belief state $b_{k}$. The selection of the specific finger index for haptic feedback proceeds in two stages: first, a state hypothesis $s_{k}$ is sampled from the current belief distribution $s_{k}\sim b_{k}$; second, the action $a_{k}$ is sampled from the policy distribution $\pi^{*}(a|s_{k})$ corresponding to the sampled state. Following the prescription of the haptic nudge by the BoMI, the system records the participant’s trial performance metrics $o_{k}=\{\texttt{RE}_{k},\texttt{SoT}_{k}\}$. The belief state is then advanced to the subsequent time step via the belief update (11). This updated belief $b_{k+1}$ encapsulates the new system knowledge and is used in the decision process for the next trial.
Figure 4: Movement Efficiency: (a) Mean SoT curves (with $95\%$ confidence intervals) of the three group participants, and (b) the mean number of trials (along with the scatter plot taken across participants) to converge to a specific SoT threshold under the three nudging policies shows fastest convergence for participants trained on POMDP-based optimal nudging policy, followed by the heuristic policy. The low p-values from two-tailed tests from the linear mixed model fits show a decrease in SoT across trials and in the number of trials required by the POMDP group for various SoT thresholds. Each scatter plot point represents the number of trials taken by a participant to achieve the SoT threshold.